Calculate the size of a rectangular sedimentation tanks to treat a flow of 12,000 m3/d. Adopt an overflow rate of 20 m3/m2-d and detention time of 4 hours. Use two units in parallel. What will be the minimum weir length required if the weir overflow rate is not to exceed 300 m3/m-d? Also check the horizontal velocity.


1
Expert's answer
2021-09-03T02:00:07-0400

"Discharge : 12,000 m^3/day\\
Overflow Rate OFR = 20 m/day\\
Depth = 3 m\\
Surface \space area \space of \space rectangular \space tank \space Surface \space area =\frac{Q}{OFR} = \frac{12000}{ 20} m^2\\
Area = 600 \\
Let \space us \space assume, length \space to \space width \space Ratio \space be 4:1\\
4x \times1x = 600\\
x^2 = \frac{600}{4}\\
x^2 = 150 m \\
X = 3. 162 m \\
4x = 12.649 m \\
Let \space us \space adope\space the \space dimension \space of ( 13 \times 4 ) m^2 \\
Detention \space time, td = \frac{D}{vs} = \frac{3}{20} day \\
td = \frac{3}{20}*24 \\
td = 3.6 hrs \\
wear \space overflow \space rate \space or \space weir \space Loading \space Rate\\
WFR : \frac{Flow \space Rate}{ Length} \\
= \frac{12000}{ 13} m^3 s^{-1} d^{-1}\\
= 312. 5 m^3/ mday"